# Equivariant Flow Matching for Symmetry-Breaking Bifurcation Problems

### v2

@@ 4 Node Graph:
For the buckling beam, we used a random walk prior similar to the approach from Brinke et al. brinke2025flow. We used an EGNN satorras2021n to predict interactions between nodes and a UNet for the time evolution.
Allen-Cahn:
@@ Appendix C Additional Results
Figure 4: Top: the results of predicting a probability distribution consisting of two Dirac deltas, using a VAE and flow matching. Bottom: the learned flow field by the trained flow matching model (contour plot), showing how samples from the Gaussian prior are pushed towards the two delta peaks (black lines).
@@
Figure 6: Comparison of predicted Allen-Cahn trajectories by the flow model with symmetric coupling with examples of possible ground truth trajectories.
Figure 7: Bifurcation diagram of the Allen-Cahn equation as the parameter $\mu$ is varied, keeping $\epsilon$ constant at 0.1. The trained model is able to model the pitchfork bifurcation.

### v3

@@ Appendix B Architectures & Training
Figure 4: Top: the results of predicting a probability distribution consisting of two Dirac deltas, using a VAE and flow matching. Middle/bottom: the learned flow field by the trained flow matching model (contour plot), showing with green lines the coupling during training between prior and targets (middle: without symmetric matching, bottom: with) and how the coupling samples from the Gaussian prior are pushed towards the two delta peaks (black lines).
@@ 4 Node Graph:
For the buckling beam, we used a random walk prior similar to the approach from Brinke et al. Brinke et al. (2025). We used an EGNN Satorras et al. (2021) to predict interactions between nodes, which internally uses UNets for the time evolution.
Allen–Cahn:
@@ Appendix C Additional Results
Figure 6: Comparison of predicted Allen–Cahn trajectories by the flow model with symmetric coupling with examples of possible ground truth trajectories.
Figure 7: Bifurcation diagram of the Allen–Cahn equation as the parameter $\mu$ is varied, keeping $\epsilon$ constant at 0.1. The trained model is able to model the pitchfork bifurcation.
Acknowledgements